OpenStudy (anonymous):

56 ÷ (-8) = a0 pleaswe help me

OpenStudy (anonymous):

If your problem was 56 ÷ 8 the answers would be 7. In your case there is just one difference, the minus sign. Your problem can be rewritten as: 56 ÷ (-8) = (-56) ÷ 8 = (-1)*(56 ÷ 8) = (-1)*7 = -7